OpenStudy (anonymous):

How do i solved by elimination 2x+5y=17 6x-5y=-9

OpenStudy (anonymous):

add them to get: 8x = 8 giving x =1. Then put x=1 into either one to get: y=3

OpenStudy (anonymous):

How do I get 8x

OpenStudy (anonymous):

add the equations: 2x+5y + 6x-5y = 17-9 8x = 8

OpenStudy (anonymous):

I dont get it. When I get the x=1 Lets say i plug it in 2(1)+5y=17. How did you get y=3

OpenStudy (anonymous):

The key is to eliminate a variable. 5x and -5x are opposites, so their sum is 0. You are left with 8x = 8. x = 1. Now 2(1) +5y =17...2 +5y =17... 5y =15..y=3

OpenStudy (anonymous):

Ok from here: 2(1)+5y=17 you get: 2+5y=17 now subtract 2 from both sides of this equation: 5y=15 finally divide both sides of this equation by giving y = 3, because 3x5=15.